OpenStudy (anonymous):

solve y+mx+b for x x+y=4 x-y=-6

OpenStudy (anonymous):

(x+y)+(x-y)=4+(-6) x+y+x-y=-2 2x=-2 x=-1 can you continue??

OpenStudy (anonymous):

no! x+y=4 -1+y=4 y=5 ok??

OpenStudy (anonymous):

get it thanks

OpenStudy (anonymous):

welcome :)